Using tabulated standard reduction potentials from
Answer #1

Reduction half cell reaction(Cathode)

Cu^{2+}(aq)+2e^-\rightarrow Cu(s) -------E^o_{(cathode)}=0.34V

Reduction half reaction(Anode)

Mg(s)\rightarrow Mg^{2+}+2e^- -------E^o_{(anode)}=-2.38V

We know,

E^o_{cell}=E^o_{cathode}-E^o_{anode}

=>E^o_{cell}=[0.34-(-2.38)]V

=>E^o_{cell}=2.72V